Without referring to a data table, rank the following compounds by standard molar entropy. . Highest S HF(g) Ne(g) F2(g) Lowest S



Which of the following reactions is likely to be product favored based on energetic considerations? (You can select more than one reaction) 2 Na(s)+ Ch(g)-> 2 NaC(s) 2 NHs(g)->No(9)+ 3 H2(9) CH4g) +2 H200)-> COg) H4 H2(9)
Answer #1

Ans 1 :

Entropy is the degree of randomness or disorderness of the system . Larger sized gaseous molecules have larger standard molar entropy and smaller sized gaseous molecules have a smaller standard molar entropy.

So here , the ranking of compounds in order of highest to lowest standard molar entropy will be :

F2 (g)

Ne (g)

HF (g)

Ans 2 : Two choices are correct :

2NH3 (g) = N2 (g) + 3H2 (g)

CH4 (g) + 2H2O (l) = CO2 (g) + 4H2 (g)

The reactions that have a negative value of gibbs free energy change are energetically favourable reactions. So if the entropy change of a reaction is positive , it is likely to be product favoured reaction.